# Mental Disorder Classification via Temporal Representation of Text

### v1

@@ Appendix B Baseline Approaches
Additionally we also use MPNet Reimers and Gurevych (2019), GPT-3.5-turbo Brown et al. (2020) and MentalLLaMa-chat-13B Yang et al. (2023a) as our baselines. They are described below:
@@ Appendix C Model Configurations and Experiments
The Feedforward network and ML models were trained on the top 30 333list of extracted features can be found here selected features extracted from time series data. Time-Series Transformer, 1D-CNN, and LSTM-based classifiers were trained on the temporal representation of data. We used the Adam optimizer, cross-entropy, as a loss function for all our experiments. In our experimental setup, we conducted training using a single NVIDIA A100-SXM4-80GB GPU. During training, we employed callbacks for model checkpointing and early stopping to optimize and prevent overfitting. We use a grid search on the validation set to search for optimal hyperparameters. Table 7 provides details of the hyperparameters used for each model in the context of Anorexia, Depression, and Self-Harm classification tasks. These hyperparameters include the learning rate (lr), batch size (# BS), and the number of epochs (# E) for each deep learning model.
@@ C.1 Additional Experiments
Method Anorexia Depression Self-Harm
lr # BS # E lr # BS # E lr # BS # E
Feedforward Network 1e-3 16 200 1e-4 32 200 1e-3 2 200
Time-Series Transformer 3e-4 16 20 3e-4 32 10 1e-3 16 20
1D-CNN 1e-3 16 50 1e-3 8 100 1e-3 16 100
LSTM 1e-2 8 50 1e-2 16 50 1e-2 8 50
Appendix D Langauge Model and Context Length
Appendix E Flos Calculation
@@ G.1 Dataset Statistics
Appendix H Cross-Domain Study: Detailed Results
As discussed in Section 9, we present the results for all six cross-domain combinations. These insights can be incorporated to design a multi-task learning framework which can help in the robust classification of low-resource mental disorders like schizophrenia, obsessive-compulsive disorder, etc.
Model Train+Val Test F1 P R
Anorexia (A)
DisorBERT A A 0.83 0.82 0.85
LSTM A A 0.79 0.84 0.74
LSTM D A 0.75 0.68 0.83
LSTM SH A 0.80 0.85 0.75
Depression (D)
DisorBERT D D 0.69 0.56 0.89
LSTM D D 0.75 0.79 0.71
LSTM A D 0.63 0.86 0.50
LSTM SH D 0.63 0.73 0.56
Self-Harm (SH)
DisorBERT SH SH 0.74 0.72 0.76
LSTM SH SH 0.83 0.93 0.75
LSTM A SH 0.78 0.85 0.72
LSTM D SH 0.69 0.65 0.77
Table 9: Results for cross-domain evaluations for all six combinations of disorders. Here, ‘A’ is anorexia, ‘D’ is depression and ‘SH’ is self-harm.

### v2

@@ Appendix B Baseline Approaches
Additionally we also use MPNet Reimers and Gurevych (2019), GPT-3.5-turbo Brown et al. (2020) and MentalLLaMa-chat-13B Yang et al. (2024) as our baselines. They are described below:
@@ Appendix C Model Configurations and Experiments
The Feedforward network and ML models were trained on the top 30 666list of extracted features can be found here selected features extracted from time series data. For feature selection, we employed the Gini impurity criterion using a Random Forest classifier, as described by Yuan et al. (2021). We selected the top 30 features based on their Gini importance scores to serve as independent variables for learning the decision space of mental disorders. Time-Series Transformer, 1D-CNN, and LSTM-based classifiers were trained on the temporal representation of data. We used the Adam optimizer, cross-entropy, as a loss function for all our experiments. In our experimental setup, we conducted training using a single NVIDIA A100-SXM4-80GB GPU. During training, we employed callbacks for model checkpointing and early stopping to optimize and prevent overfitting. We use a grid search on the validation set to search for optimal hyperparameters. Table 8 provides details of the hyperparameters used for each model in the context of Anorexia, Depression, and Self-Harm classification tasks. These hyperparameters include the learning rate (lr), batch size (# BS), and the number of epochs (# E) for each deep learning model.
@@ C.1 Additional Experiments
Method Anorexia Depression Self-Harm
lr # BS # E lr # BS # E lr # BS # E
Feedforward Network 1e-3 16 200 1e-4 32 200 1e-3 2 200
Time-Series Transformer 3e-4 16 20 3e-4 32 10 1e-3 16 20
1D-CNN 1e-3 16 50 1e-3 8 100 1e-3 16 100
LSTM 1e-2 8 50 1e-2 16 50 1e-2 8 50
Appendix D Langauge Model and Context Length
Appendix E FLOPs Calculation